OpenStudy (anonymous):

Is the function even or odd? y= 1/2cos2x*sin(3Pi/2 - 2x)+3

OpenStudy (anonymous):

what is your answer?

ganeshie8 (ganeshie8):

before testing f(-x), simplify the function a bit : y= 1/2cos2x*sin(3Pi/2 - 2x)+3 y= 1/2cos2x*sin(pi + pi/2 -2x) + 3 y= 1/2cos2x*-sin(pi/2 -2x) + 3 y= 1/2cos2x*-cos(2x) + 3 y= -1/2cos^2(2x) + 3 now test to see if f(x) = f(-x)

OpenStudy (anonymous):

IF f(-x)=f(x),function is even and if f(-x)=-f(x) ,the function is termed as odd function

ganeshie8 (ganeshie8):

are we getting f(x) = -f(x) ? our simplified y=f(x) is this : y = f(x) = -1/2cos^2(2x) + 3 lets find f(-x) and see wat u get

ganeshie8 (ganeshie8):

f(x) = -1/2cos^2(2x) + 3 ---------(1) f(-x) = -1/2cos^2(-2x) + 3 = -1/2cos^2(2x) + 3 -------(2)

ganeshie8 (ganeshie8):

(1) = (2), so... ?

OpenStudy (anonymous):

The function is even

OpenStudy (anonymous):

Thanks a lot!

ganeshie8 (ganeshie8):

wonderful !

ganeshie8 (ganeshie8):

yw